# Vision-Language Models for Medical Report Generation and Visual Question Answering: A Review

### v1

@@ 3.2.4 Fine-Tuning Techniques
Following the training of VLMs, a common practice involves fine-tuning them on smaller datasets tailored to specific downstream tasks. One fine-tuning approach in the context of ITM pre-training is known as supervised fine-tuning (SFT). In the initial phase, the model undergoes pre-training on an extensive image-text dataset, establishing a foundational understanding of the complex interplay between visual and textual components. Subsequently, meticulous fine-tuning takes place on a more focused dataset, curated to match the nuances of the targeted application. This dual-phase strategy, encompassing broad pre-training and task-specific fine-tuning, enables the model to benefit from large-scale generalization while seamlessly adapting to the intricacies of particular applications.
In contrast, reinforcement learning from human feedback (RLHF) is a distinct fine-tuning approach employed to enhance VLMs through a systematic sequence of steps. Initially, an existing VLM serves as the foundation, with the option for potential fine-tuning using additional textual data. Subsequently, a detailed reward model is constructed by ranking text outputs generated by the VLM based on human preferences. The fine-tuning phase employs a policy-gradient reinforcement learning algorithm, such as Proximal Policy Optimization (PPO), to adjust the VLM’s parameters. The reward function combines scalar rewards from the trained model with a penalty mechanism that discourages substantial deviations from the pre-trained model. This iterative process ensures the optimization of the VLM, tailoring it to specific visual and textual tasks, aligning it with human preferences, and thereby enhancing its overall performance and adaptability.
@@
Instruction fine-tuning refers to the process of refining a pre-trained language model by providing specific instructions or guidance tailored to a particular task or application. This process typically involves exposing the model to examples or prompts related to the desired instructions and updating its parameters based on the feedback received during this task-specific training phase. RaDialog [Pel+23] employs this fine-tuning technique.
3.3 Adapter Training
In this section, we explore strategies for adapting VLMs while keeping the model’s parameters frozen and only updating newly added layers.
@@ 3.3 Adapter Training
In recent years, parameter-efficient fine-tuning (PEFT) has gained prominence, encompassing various techniques and strategies that aim to make the most effective use of parameters during the fine-tuning process, particularly in scenarios with limited labeled data for the target task. One strategy is incorporating task-specific parameters, known as adapters, into a pre-trained model while retaining its original parameters. The architecture of adapter modules typically incorporates a bottleneck structure, projecting original features into a reduced dimension, applying non-linearity, and then projecting back to the original dimension. This thoughtful design ensures parameter efficiency by limiting the number of added parameters per task. Integrated after each layer of the pre-trained model, adapter modules capture task-specific details while preserving shared parameters, enabling the model’s seamless extension to new tasks without significant interference with previously acquired knowledge.
@@
The common adapter-based method is low-rank adaptation (LoRA) [Hu+22]. In LoRA, the adaptation process involves fine-tuning two smaller matrices that approximate the larger weight matrix of the pre-trained language model. These smaller matrices constitute the LoRA adapter modules, and the approach focuses on making low-rank modifications to efficiently adapt the model for specific tasks. Pre-trained LLMs that are part of medical VLMs architecture are often fine-tuned using LoRA (e.g., Visual Med-Alpaca [Shu+23] and RaDialog [Pel+23]).
@@
Prompt tuning [LAC21] involves creating continuous vector representations as input hints. This allows the model to dynamically generate effective prompts during training. Continuous refinement of prompts significantly enhances the model’s ability to generate accurate and contextually relevant responses. This iterative process allows the model to adapt its behavior based on an evolving understanding of the task. Prompt tuning was used by VLMs like Qwen-VL [Bai+23] and InstructBLIP [Dai+23].
@@
Prefix token tuning [LL21] adds task-specific vectors to the input, specifically to the initial tokens known as prefix tokens, to guide the model’s behavior for a given task. For instance, VL-T5 [Cho+21] utilized different prefixes for questions from various datasets. These vectors can be trained and updated independently while keeping the remaining pre-trained model parameters frozen. Prefix token tuning allows for task-specific adaptation without compromising the pre-trained knowledge encoded in the majority of the model’s parameters.
3.4 Prompt Engineering
@@
In this section, we explore strategies for adapting VLMs while keeping the model’s parameters frozen and refraining from any updates.
@@
Prompt engineering [Gu+23] is a technique that involves enhancing a large pre-trained model with task-specific instructions, referred to as prompts, to tailor the model for specific tasks. Examples include instructing the model to generate a radiology report for a specific image (e.g., RAMM [Pel+23]). Prompt engineering can also expose the VLM to a sequence of interconnected examples or prompts, guiding it to a desired output. Another approach is to incorporate progressively structured instructions or questions, refining focus and enhancing the model’s ability to generate coherent and contextually relevant responses.
@@
Retrieval augmented generation (RAG) is a form of prompt engineering that involves strategically crafting prompts for both retrieval and generation phases, allowing for an adaptive and efficient process that leverages external knowledge sources to enhance generative tasks. While the original concept of RAG was developed in the context of NLP [Lew+20], the principles behind retrieval and generation can be extended to multimodal learning [Zha+23c], including VLMs. RAG has been used in vision-language tasks like VQA (e.g., RAMM [Yua+23]) and RG (e.g., CXR-RePaiR-Gen [Ran+23a]). RAG begins with a retrieval component, which is usually a pre-trained model designed for information retrieval. This versatile component excels in extracting pertinent information from extensive datasets, catering to various modalities such as images, text, codes, video, or audio when presented with diverse inputs [Zha+23c]. Following the retrieval phase, the model returns a set of contexts related to the given input. The second component is a generative language model, often based on architectures like GPT. This component takes the input and the retrieved context and generates the final output. The generated output is conditioned not only on the input but also on the information extracted from the retrieved context. An intrinsic advantage of RAG lies in its capacity to reduce the reliance on extensive labeled datasets. By tapping into external knowledge sources, the model becomes more data-efficient and adept at handling a spectrum of tasks. While the base model is typically frozen during RAG, there are instances, as seen in RAMM [Yua+23], where model parameters are updated in the process.
@@ 3.5 Downstream Tasks
Multimodal downstream tasks leverage the acquired knowledge from pre-training VLMs to excel in diverse applications that require a joint understanding of visual and textual data. A prominent example is report generation (RG) task, which centers on creating a comprehensive summary report of visual data. In the medical domain, RG plays a crucial role in automatically summarizing diagnostic imaging results and reducing the workload of report writing [TLZ23]. For instance, in radiology, a report generation system could analyze a set of medical images such as X-rays, CT scans, or MRIs, and generate a detailed report summarizing the observed abnormalities, their locations, and potential implications for diagnosis or treatment [LTS23]. A radiology report usually has several sections: Examination (type of exam), Indication (reasons for the examination), Comparison (prior exams), Technique (scanning method) Findings (detailed observations made by a radiologist), and Impression (summary of the major findings) [MHC20]. In the context of RG, VLMs are usually designed to generate Findings and Impression sections [Tha+23]. Currently, VLMs tailored for RG are predominantly utilized for radiology images, with lesser application in other medical imaging domains such as histopathology [SB23], robotic surgery [Xu+21], and ophthalmology [Li+22].
@@
Another significant visual-language understanding task is visual question answering (VQA), where the model needs to comprehend images or videos and the posed question to provide a relevant and accurate response [Ant+15]. The spectrum of questions encountered in VQA is broad, encompassing inquiries about the presence of specific objects, their locations, or distinctive properties within the image. In a medical context [Lin+23a], this may involve questions regarding the presence of medical conditions or abnormalities, such as “What abnormality is seen in the image?” [Ion+21] or “Is there gastric fullness?” [Lau+18]. Other queries may delve into details like the imaging method used [Aba+19], the organ system involved [Lau+18], or the presence of specific anatomical structures [Liu+21].
@@
Questions in VQA fall into two categories. Open-ended questions elicit responses in the form of phrases or sentences, fostering detailed and nuanced answers [Tha+23]. On the other hand, closed-ended questions are designed to prompt limited responses, often with predetermined options, such as a short list of multiple choices, a yes/no response, or a numeric rating [Baz+23]. The task of VQA is commonly approached as either a classification task or a generation task or both [Lin+23a]. In the classification approach, models select the correct answer from a predefined set, while in the generation task, models produce free-form textual responses unconstrained by predefined options.
@@
Beyond VQA and RG, a spectrum of tasks contributes to the rich landscape of vision-language understanding [Che+23]. For instance, referring expression comprehension entails a model locating the specific area or object in an image that the given phrase or sentence refers to [ZNC18]. Visual commonsense reasoning involves answering questions about an image, typically presented in a multiple-choice format, and justifying the answer based on the model’s understanding of the image and common sense knowledge [Zel+19]. Vision-language retrieval focuses on either generating or retrieving relevant information from images using textual data, or vice versa, obtaining information from text using visual data [Zhe+19]. In the context of visual captioning, the model’s role is to generate a concise, text-based description of either an image [SDK23]. It is worth highlighting that some of these tasks can seamlessly transition from images to videos showcasing the adaptability and versatility of VLMs across diverse visual contexts [Gan+22].
@@
[Kay+22] evidence labels
CXR-PRO $374,139$ reports, PN
[RCR22] $374,139$ radiographs
@@
[Boe+22]
IU-Xray or Open-I $7,470$ labels Openi
[Dem+15]
MedICaT $217,060$, annotations; inline GH
[Dem+15] $7,507$ references to ROCO figures
@@
[Lin+23]
SLAKE $14,028$ $642$ annotated images, MedVQA
[Liu+21] $5,232$ medical triplets
VQA-RAD $3,515$ $315$ radiology images Osf
[Lau+18]
@@
[Aba+20] images and questions for VQG
@@
[Ion+21] images and questions for VQG
@@
[All+19] 97 frames
EndoVis 2018 $11,783$ bounding box annotations; GH
[All+20] 2,007 frames challenge
Table 1: A list of medical vision-language datasets.
4 Medical Vision-Language Datasets
@@
The adaptation of VLMs to the healthcare sector is achieved through their pre-training and fine-tuning using specialized medical datasets. Below is the list of vision-language datasets available to the public that contain medical image-text pairs or question-answer (QA) pairs. Most of them are employed by medical VLMs described in Section 6 for pre-training, fine-tuning, and evaluation of VQA and RG tasks. The comparative analysis of these datasets is presented in Table 1.
@@
  • ROCO (Radiology Objects in Context) [Pel+18] is a dataset composed of image-caption pairs extracted from the open-access biomedical literature database PubMed Central (PMC). ROCO is stratified into two categories: radiology and out-of-class. The radiology group includes $81,825$ radiology images, including computer tomography (CT), ultrasound, x-ray, fluoroscopy, positron emission tomography (PET), mammography, magnetic resonance imaging (MRI), angiography, and PET-CT. The out-of-class group has $6,127$ images, including synthetic radiology images, clinical photos, portraits, compound radiology images, and digital art. Each image is accompanied by a corresponding caption, keyword, Unified Medical Language System (UMLS) semantic types (SemTypes), UMLS concept unique identifiers (CUIs), and a download link. To facilitate model training, the dataset is randomly split into a training set ($65,460$ radiology and $4,902$ out-of-class images), a validation set ($8,183$ radiology and $612$ out-of-class images), and a test set ($8,182$ radiology and $613$ out-of-class images) using an 80/10/10 split ratio, respectively.
@@
  • MIMIC-CXR (Medical Information Mart for Intensive Care - Chest X-Ray) [Joh+19a] collection encompasses $377,110$ chest X-rays paired with $227,835$ associated free-text radiology reports. The dataset is derived from de-identified radiographic studies conducted at the Beth Israel Deaconess Medical Center in Boston, MA. Each imaging study within the MIMIC-CXR dataset consists of one or more images, typically featuring lateral and from back-to-front (posteroanterior, PA) views in Digital Imaging and Communications in Medicine (DICOM) format.

### v2

@@ 3.2.5 Fine-Tuning Techniques
Following the training, a common practice involves fine-tuning VLMs on smaller datasets tailored to specific downstream tasks.
Supervised Fine-Tuning (SFT)
@@
Before employing SFT, the VLM is pre-training on an extensive image-text dataset, establishing a foundational understanding of the complex relationship between visual and textual representations. SFT involves meticulous fine-tuning on a more focused dataset, curated to match the nuances of the targeted application. This dual-phase strategy, encompassing broad pre-training and task-specific fine-tuning, enables the model to benefit from large-scale generalization while seamlessly adapting to the intricacies of particular applications [ouyang2022training].
Reinforcement Learning from Human Feedback (RLHF)
RLHF is a distinct fine-tuning approach employed to enhance VLMs through the incorporation of human preferences during fine-tuning [ouyang2022training, lambert2022illustrating, ziegler2020finetuning]. RLHF initiates with an initial model, incorporating human-generated rankings of its outputs to construct a detailed reward model. In contrast to traditional reinforcement learning (RL) [sutton1998rlintro, coronato2020rl], which relies solely on environmental interactions, RLHF strategically integrates human feedback. This human-in-the-loop approach provides a more nuanced and expert-informed methodology, allowing for the fine-tuning of VLMs in alignment with human preferences, ultimately leading to improved model outcomes.
Instruction Fine-Tuning (IFT)
@@
IFT refers to the process of refining a pre-trained language model by providing specific instructions or guidance tailored to a particular task or application [ren2024learning-IFT]. This process typically involves exposing the model to examples or prompts related to the desired instructions and updating its parameters based on the feedback received during this task-specific training phase. Medical VLM, RaDialog [pellegrini2023radialog], employs this fine-tuning technique.
3.3 Parameter-Efficient Fine-Tuning (PEFT)
@@
In this section, we explore strategies for adapting VLMs while keeping the model’s parameters frozen and only updating newly added layers. In recent years, PEFT has gained prominence, encompassing various techniques and strategies that aim to make the most effective use of parameters during the fine-tuning process, particularly in scenarios with limited labeled data for the target task. The main strategy of PEFT is incorporating task-specific parameters, known as adapters, into a pre-trained model while retaining its original parameters. The architecture of adapter modules typically incorporates a bottleneck structure, projecting original features into a reduced dimension, applying non-linearity, and then projecting back to the original dimension. This thoughtful design ensures parameter efficiency by limiting the number of added parameters per task. Integrated after each layer of the pre-trained model, adapter modules capture task-specific details while preserving shared parameters, enabling the model’s seamless extension to new tasks without significant interference with previously acquired knowledge.
3.3.1 Low-Rank Adaptation (LoRA)
@@
LoRA is a common adapter-based method is [hu2022lora]. The adaptation process involves fine-tuning two smaller low-rank matrices that are decompositions of the larger weight matrix of the pre-trained model. These smaller matrices constitute the LoRA adapter modules, and the approach focuses on making low-rank modifications to adapt the model for specific tasks efficiently. Pre-trained LLMs that are part of medical VLMs architecture are often fine-tuned using LoRA (e.g., Visual Med-Alpaca [shu2023vismedalpaca] and RaDialog [pellegrini2023radialog]).
3.3.2 Prompt Tuning
@@
Prompt tuning involves creating continuous vector representations as input hints [lester2021prompttuning]. This allows the model to generate effective prompts during training dynamically. Continuous refinement of prompts significantly enhances the model’s ability to generate accurate and contextually relevant responses. This iterative process allows the model to adapt its behavior based on an evolving understanding of the task. VLMs like Qwen-VL and InstructBLIP used prompt tuning [bai2023qwenvl, dai2023instructblip].
3.3.3 Prefix Token Tuning
@@
Prefix token tuning adds task-specific vectors to the input, specifically to the initial tokens known as prefix tokens, to guide the model’s behavior for a given task [li2021prefixtuning]. For instance, VL-T5 utilized different prefixes for questions from various datasets [cho2021vlt5] . These vectors can be trained and updated independently while keeping the remaining pre-trained model parameters frozen. Prefix token tuning allows for task-specific adaptation without compromising the pre-trained knowledge encoded in the majority of the model’s parameters.
3.4 In-Context Learning
@@
In this section, we explore strategies for adapting VLMs using the context only, keeping the model’s parameters (and PEFT/LoRA adapters, if any) frozen. In our settings, in-context learning may be considered as using LLMs or VLMs for inference only.
3.4.1 Prompt Engineering
@@
Prompt engineering is a technique that involves enhancing a large pre-trained model with task-specific instructions, referred to as prompts, to tailor the model’s output for specific tasks [gu2023promptvlm]. Examples include instructing the model to generate a radiology report for a specific image (e.g., RAMM [pellegrini2023radialog]). Prompt engineering can also expose the VLM to a sequence of interconnected examples or prompts, guiding it to a desired output. Another approach incorporates progressively structured instructions or questions, refining focus and enhancing the model’s ability to generate coherent and contextually relevant responses [gu2023promptvlm].
3.4.2 Retrieval augmented generation (RAG)
@@
RAG is a form of prompt engineering that involves strategically crafting prompts for both retrieval and generation phases, allowing for an adaptive and efficient process that leverages external knowledge sources to enhance generative tasks. While the original concept of RAG was developed in the context of NLP [lewis2020rag], the principles behind retrieval and generation can be extended to multimodal learning [zhao2023multimodalrag], including VLMs. RAG has been used in medical VLMs for tasks like VQA (e.g., RAMM [yuan2023ramm]) and RG (e.g., CXR-RePaiR-Gen [ranjit2023cxrrepairgen]). RAG begins with a retrieval component, which is usually a pre-trained model designed for information retrieval. This versatile component excels in extracting pertinent information from extensive datasets, catering to various modalities such as images, text, codes, video, or audio when presented with diverse inputs [zhao2023multimodalrag]. Following the retrieval phase, the model returns a set of contexts related to the given input. The second component is a generative LLM. This component takes the input and the retrieved context and generates the final output. The generated output is conditioned not only on the input but also on the information extracted from the retrieved context. An intrinsic advantage of RAG lies in its capacity to reduce the reliance on extensive labeled datasets. While the base model is typically frozen during RAG, there are instances, as seen in RAMM [yuan2023ramm], where model parameters are updated in the process.
@@ 3.5 Downstream Tasks
Multimodal downstream tasks leverage the acquired knowledge from pre-training VLMs to excel in diverse applications that require a joint understanding of visual and textual data.
3.5.1 Report Generation (RG)
@@
RG is a prominent example of a typical medical VLM task, which centers on creating a comprehensive summary report of visual data. RG plays a crucial role in automatically summarizing diagnostic imaging results and reducing the workload of report writing [monshi2020radreportsurvey, ting2023mrgsurvey, Mohsan2023]. For instance, in radiology, a report generation system could analyze a set of medical images such as X-rays, CT scans, or MRIs and generate a detailed report summarizing the observed abnormalities, their locations, and potential implications for diagnosis or treatment [liu2023rrgreview]. A radiology report usually has several sections: (1) Examination (type of exam), (2) Indication (reasons for the examination), (3) Comparison (prior exams), (4) Technique (scanning method) (5) Findings (detailed observations made by a radiologist), and (6) Impression (summary of the major findings) [mabotuwana2020radreportfram]. In the context of RG, VLMs are usually designed to generate Findings and Impression sections [thawkar2023xraygpt]. Currently, VLMs tailored for RG are predominantly utilized for radiology images, with lesser application in other medical imaging domains such as pathology [sengupta2023pathologyrg], robotic surgery [Xu2021rginsurgery], and ophthalmology [li2022rginophthalmology].
3.5.2 Visual Question Answering (VQA)
@@
VQA is another important visual-language understanding task, where the model needs to comprehend images or videos and the posed question to provide a relevant and accurate response [antol2015vqa]. The spectrum of questions encountered in VQA is broad, encompassing inquiries about the presence of specific objects, their locations, or distinctive properties within the image. In the medical context [lin2023medvqa], this may involve questions regarding the presence of medical conditions or abnormalities, such as “What abnormality is seen in the image?” [ionescu2021vqamed21] or “Is there gastric fullness?” [lau2018vqarad]. Other queries may delve into details like the imaging method used [abacha2019vqamed19], the organ system involved [lau2018vqarad], or the presence of specific anatomical structures [liu2021slake].
@@
Questions in VQA fall into two categories. Open-ended questions elicit responses in the form of phrases or sentences, fostering detailed and nuanced answers [thawkar2023xraygpt]. On the other hand, closed-ended questions are designed to prompt limited responses, often with predetermined options, such as a short list of multiple choices, a yes/no response, or a numeric rating [bazi2023vqamedim]. The task of VQA is commonly approached as either a classification task, a generation task, or both [lin2023medvqa]. In the classification approach, models select the correct answer from a predefined set, while in the generation task, models produce free-form textual responses unconstrained by predefined options.
3.5.3 Other Tasks
@@
Beyond VQA and RG, a spectrum of VLM tasks exist for the vision-language understanding [chen2022vlp]. For instance, referring expression comprehension entails a model locating the specific area or object in an image that the given phrase or sentence refers to [zhang2018gre]. Visual commonsense reasoning involves answering questions about an image, typically presented in a multiple-choice format, and justifying the answer based on the model’s understanding of the image and common sense knowledge [zellers2019vcr]. Vision-language retrieval focuses on either generating or retrieving relevant information from images using textual data, or vice versa, obtaining information from text using visual data [zhen2019dscretrival]. In the context of visual captioning, the model’s role is to generate a concise, text-based description of either an image [sharma2023captioning]. It is worth highlighting that some of these tasks can seamlessly transition from images to videos, showcasing the adaptability and versatility of VLMs across diverse visual contexts [gan2022vlmsurvey].
Table 1: A list of datasets used for developing medical VLMs.
@@
[kayser2022mimicnle] evidence labels
CXR-PRO $374,139$ radiographs and PN
[ramesh2022cxrredone] $374,139$ reports but not paired
@@
[boecking2022mscxr]
IU-Xray or Open-I $7,470$ labels Site
[demnerfushman2015openi]
MedICaT $224,567$ annotations; inline GH
[subramanian2020medicat] references to ROCO figures
@@
[lin2023pmcclip]
SLAKE $14,028$ $642$ annotated images, Site
[liu2021slake] $5,232$ medical triplets
VQA-RAD $3,515$ $315$ radiology images Site
[lau2018vqarad]
@@
[abacha2020vqamed20] images and questions for VQG
@@
[ionescu2021vqamed21] images and questions for VQG
@@
[allan2019endovis17] 97 frames
EndoVis 2018 $11,783$ bounding box annotations; GH + Site
[allan2020endovis18] 2,007 frames
Note: Abbreviations used are: GH - GitHub, HF - Hugging Face, and PN - PhysioNet
4 Medical VLMs
4.1 Medical Datasets for VLMs
@@
The adaptation of VLMs to various medical tasks is achieved through their pre-training and fine-tuning using specialized task-specific datasets. Below is the list of vision-language datasets available in the public domain that contain medical image-text pairs or question-answer (QA) pairs. Most of them are employed by medical VLMs described in Section 4.3 for pre-training, fine-tuning, and evaluating VQA and RG tasks. The comparative analysis of these datasets is presented in Table 1. The last column in Table 1 provides a link to the source of the data on the web with the following abbreviations: GH - GitHub, PN - PhysioNet, and HF - Hugging Face.
4.1.1 Radiology Objects in Context (ROCO)
@@
ROCO is a dataset composed of image-caption pairs extracted from the open-access biomedical literature database PubMed Central (PMC) [pelka2018roco]. ROCO is stratified into two categories: radiology and out-of-class. The radiology group includes $81,825$ radiology images, including computer tomography (CT), ultrasound, x-ray, fluoroscopy, positron emission tomography (PET), mammography, magnetic resonance imaging (MRI), angiography, and PET-CT. The out-of-class group has $6,127$ images, including synthetic radiology images, clinical photos, portraits, compound radiology images, and digital art. Each image is accompanied by a corresponding caption, keyword, Unified Medical Language System (UMLS) semantic types (SemTypes), UMLS concept unique identifiers (CUIs), and a download link. To facilitate model training, the dataset is randomly split into a training set ($65,460$ radiology and $4,902$ out-of-class images), a validation set ($8,183$ radiology and $612$ out-of-class images), and a test set ($8,182$ radiology and $613$ out-of-class images) using an 80/10/10 split ratio, respectively.
4.1.2 Medical Information Mart for Intensive Care - Chest X-Ray (MIMIC-CXR)
@@
MIMIC-CXR collection encompasses $377,110$ chest X-rays paired with $227,835$ associated free-text radiology reports [johnson2019mimiccxr]. The dataset is derived from de-identified radiographic studies conducted at the Beth Israel Deaconess Medical Center in Boston, MA. Each imaging study within the MIMIC-CXR dataset consists of one or more images, typically featuring lateral and from back-to-front (posteroanterior, PA) views in Digital Imaging and Communications in Medicine (DICOM) format.
4.1.3 MIMIC-CXR-JPG